40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Nov 20 14:48:55 2011, kiwamu, Update, General, recovery from the power shutdown 
    Reply  Sun Nov 20 15:33:37 2011, rana, Update, General, recovery from the power shutdown 
    Reply  Wed Nov 23 22:58:33 2011, Koji, Update, General, c1iscaux2 is back (re: recovery from the power shutdown) 
    Reply  Thu Nov 24 01:28:09 2011, kiwamu, Update, General, 1X8 sorensen came back 
Message ID: 5963     Entry time: Sun Nov 20 14:48:55 2011     Reply to this: 5965   5992   5993
Author: kiwamu 
Type: Update 
Category: General 
Subject: recovery from the power shutdown 

Recovery from the power shutdown

 - Turned on the raid disk of linux1.

 - Woke linux1 up. No fsck this time.

 - Woke up all the lab machines.

 - Turned on all the electronics racks' AC powers

 - Woke up fb and then front end machine (the raid for fb had been already up as I turned on the AC powers)

 - Turned on all the electronics racks' DC powers (Sorensens, Kepcos, and etc.)

 - Turned on the Marcnois which is driving the RF generation box.

 - Woke up all the lasers (PSL and End lasers)

 - Some burtrestoring (c1ioo, c1sus, c1susaux, c1msc, c1psl, c1iool0, c1auxey, c1auxex, c1oaf, c1pem)

 - Ran autolockMC scripts on op340m => After relocking of PMC a lock of MC was acquired immediately.

 - Turned on the PZT HV drivers.


Some issues

 - One of the Sorensens in 1X8 rack is showing the current limit sign. This is exactly the same situation as we saw before (#5592).

       Currently it's off. It needs an investigation to find who is drawing such a large amount of current.

 - C1SCX is not properly running. Rebooting the machine didn't help. This needs to be fixed.

       The symptom is : (1) all the values are frozen in the screens. (2) the c1scx status screens shows NO SYNC sign. (3) however the timing board looks blinking happily.

 - One of the VME rack on 1X3 is not showing the +/-15V green LED lights.

   This is the one on very upper side of the rack, which contains the old c1lsc machine and c1iscaux2. If we are still using c1iscaux2, it needs to be fixed.

ELOG V3.1.3-